OpenStudy (anonymous):

f(x) = x3 + 3x2 – 10x – 24. I need to identify the correct factors of f(x). These are suggestions: (x – 1) (x + 2) (x – 3) (x + 4) (x + 6) (x – 12) List the correct factors. Then justify your selections with complete sentences.

OpenStudy (anonymous):

@phi @satellite73 @Hero

hero (hero):

x - 1 is a factor of f(x) if f(1) = 0 x + 2 is a factor of f(x) if f(-2) = 0 x - 3 is a factor of f(x) if f(3) = 0 See a pattern?

OpenStudy (anonymous):

Ohh So i solve for them until one get me an answer of 0?

OpenStudy (anonymous):

It will probably be more than one

hero (hero):

Now to check if x - 1 is a factor: f(1) = (1)^3 + 3(1)^2 - 10(1) - 24 = ?

hero (hero):

Yes, when you find all of the factors that equal zero, then you will have all of them.

OpenStudy (anonymous):

Wouldn't it be (f-1) = (-1)^3 + 3(-1)^2 - 10(-1) - 24 = ?

hero (hero):

No, don't confuse yourself To find the factors, this is what you do: x - 1 = 0 x = 1

hero (hero):

Which means since x = 1 is a possible factor, we have to substitute 1 in place of x for f(x). So find f(1). If it equals zero, then you have found one of the factors of f(x). If it doesn't then move on to the next possible factor.

OpenStudy (anonymous):

So every factor would be positive?

hero (hero):

Think back to the zero product property in order to understand

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

I have another question, think you could help? • 2 is a zero of g(x) because long division with (x + 2) results in a remainder of 0. • 2 is a zero of g(x) because g(2) = 0. • 2 is a zero of g(x) because synthetic division with 2 results in a remainder of 0. Correct the reasoning of any inaccurate statements. Make sure you reference any theorems that support your justifications.

hero (hero):

This site was designed so that one question per thread would be posted. Notice that in the blue box at the top of this thread is reserved for questions. The white boxes are reserved for responses.

hero (hero):

In other words, if you have another question, the correct course of action is to close this question, then post a new question.

OpenStudy (anonymous):

If I start a new thread and tag you could you help? I figured that since you were already here I might as well post it...

hero (hero):

If you have another question, close this one, and post the other question. If you want to tag me or someone else, that's fine.